OpenStudy (anonymous):

solve the indefinite integral fnInt(xcot(10x^2))dx

OpenStudy (anonymous):

Substitute \(u=10x^2\), then \(\dfrac{1}{20}du=x~dx\): \[\int x\cot10x^2~dx=\frac{1}{20}\int\cot u~du\]